### v1

# Market Reaction to News Flows in Supply Chain Networks

Abstract
This study examines whether positive news about firms increases their stock prices and, moreover, whether it increases stock prices of the firms’ suppliers and customers, using a large sample of publicly listed firms across the world and another of Japanese listed firms. The level of positiveness of each news article is determined by FinBERT, a natural language processing model fine-tuned specifically for financial information. Supply chains of firms across the world are identified mostly by financial statements, while those of Japanese firms are taken from large-scale firm-level surveys. We find that positive news increases the change rate of stock prices of firms mentioned in the news before its disclosure, most likely because of diffusion of information through informal channels. Positive news also raises stock prices of the firms’ suppliers and customers before its disclosure, confirming propagation of market values through supply chains. In addition, we generally find a larger post-news effect on stock prices of the mentioned firms and their suppliers and customers than the pre-news effect. The positive difference between the post- and pre-news effects can be considered as the net effect of the disclosure of positive news, controlling for informal information diffusion. However, the post-news effect on suppliers and customers in Japan is smaller than the pre-news effect, a result opposite to those from firms across the world. This notable result is possibly because supply chain links of Japanese firms are stronger than global supply chains while such knowledge is restricted to selected investors.
Introduction
Firms interact with and influence each other through various types of networks. Major firm networks that are found to play a significant role in firms’ behaviors and performance in the literature include financial networks through ownership [1, 2, 3] and knowledge networks through research collaboration.[4, 5, 6, 7, 8] Another type of network recently paid attention in the academic literature and business practices is supply chains formed through transactions of materials, parts, and components.[9, 10] One reason for the growing attention is that supply chains have expanded globally, linking firms in the world within a small number of steps with each other. [11]
Global supply chains have rapidly expanded because interconnected firms through supply chains can greatly benefit from the network. Most notably, supply chains can empower firms by enhancing efficiency of production processes and thus optimizing firms’ economic performance.[12, 13] For instance, by procuring and manufacturing different goods in various locations based on each location’s comparative advantages and cost effectiveness, firms can streamline their operations as a network, ultimately leading to reductions in production costs and improvements in productivity.[14] Moreover, it is often evidenced that productivity spills over through supply chains from upstream suppliers to downstream client firms because of high-quality materials, parts, components, and services and from downstream to upstream firms because of learning of knowledge and technology.[15, 7, 16]
Conversely, supply chains can be a channel of negative shocks that aggravate firm performance, particularly once supply chains are disrupted due to unexpected events, such as natural disasters and geopolitical instability.[17, 18, 19] Such negative shocks examined in the literature include the Great East Japan Earthquake in 2011,[20] the massive flood in Thailand in 2011,[21] the COVID-19 pandemic from 2020 to 2022,[22], and the Russia-Ukraine war that started in 2022.[23]
Therefore, positive or negative shocks propagate through supply chains, affecting performance of firms linked through supply chains. Because the corporate value of publicly listed firms is evaluated in the stock market, [24, 25, 26] a shock to a firm is expected to affect not only its own stock price but also the stock price of its suppliers and customer firms. For example, when Boeing 737 MAX crashed in 2019 and 2020, the stock price of Boeing, its supplier, such as GE and Allegheny Technologies, and its customers, such as American Airlines and Southwest Airlines, dropped sharply.[27, 28, 29, 30, 31]. Also, a misconduct by Daihatsu Motor Co., a Japanese automobile manufacturer affiliated with Toyota, in 2023 negatively affected the stock price of Daihatsu, Toyota, and their suppliers and customers, such as Aisin.[32, 33]
These examples show that the shock can propagate through supply chains and affect stock prices of firms linked through supply chains. Although previous studies have examined the direct impact of shocks to firms on their own stock prices[34, 35, 36], no study has investigated the diffusion of shocks to other firms’ stock prices through supply chains to the best of our knowledge. To fill the gap, this study examines this possible diffusion, using two large firm-level samples, one for firms across the world and the other for Japanese firms. Specifically, we test whether the level of negative sentiment about a firm in a news article has a positive impact on the change rate of the stock price of the firm and its suppliers and customers before and after the disclosure of the news.
We contribute to the literature by utilizing large-scale data in many respects. First, our firm-level data covers all publicly listed firms in the world and in Japan and identify detailed supply chains among them using financial statements, news articles, websites, and firm-level surveys. Second, our data on news articles are taken from the NewsArchive of Thomson Reuters and cover more than 20 million articles. Third, we employ daily stock prices of all listed firms to examine both short- and long-run effects of positive news. Another notable contribution of this study is that we utilize sentiment analysis to determine the positiveness of each news article. Prior research on sentiment analysis has predominantly focused on neural networks [37, 38], and recently, Bidirectional Encoder Representations from Transformers (BERT)[39] has garnered the most acclaim as the state-of-the-art approach [35]. We take advantage of FinBERT,[40] an extended model of BERT particularly for financial information.
Data
This study utilizes four major data sources. First, our information on news relies on news articles written in English included in the NewsArchive provided by Thomson Reuters for the period 2003-2016. The information in each news article includes the date, headline, main text, International Securities Identification Numbers (ISIN) of firms mentioned in the article, and language. As we will explain later in detail, our sentiment analysis to determine whether each news article is positive or negative about firms mentioned there utilizes the main text. The average number of words in an article is 243, whereas its median is 136 and maximum is 3911. Although the total number of news articles in the NewsArchve for the period examined is 20,803,561, we focus on 3,447,425 news articles that mention any firm publicly listed in 105 stock markets in the world. 90% of these news articles mention 3 or fewer firms, whereas the maximum number of mentioned firms in an article is 33 (Figure 1). The number of news articles that mention a particular firm during the period 2003-2016 is quite skewed. **% of firms were never mentioned, and **% were mentioned only once, while 0.1% were mentioned more than 10,000 times (Figure 2).
Second, we utilize daily stock prices of currently listed firms in stock markets in the world taken from the Eikon database provided by Refinitiv, one of the largest providers of financial markets data and infrastructure, using its Application Programming Interface (API). We drop firms listed in the past but not currently from our sample, because their stock prices are not available using the API. We particularly define the daily stock price of each firm as its closing price. In addition to stock prices of each firm, we utilize the Refinitiv index for each stock market that indicates the average of stock prices in the market to control for the market-specific time trend. The Refinitiv index is available for 51 major markets that cover 92% of listed firms worldwide.
Third, to identify global supply chains among publicly listed firms in the world, we rely on data from FactSet that cover a number of firms across the world for the period 2003-2016 obtained as of January 18, 2023. The FactSet data include major clients of each firm reported in its financial statement. The US, Japanese, and many other governments require every listed firm to disclose its major clients with which the sales are more than 10% of its total sales, at least in the period examined in this study [41]. Therefore, based on published financial statements and, in addition, supplemental information from news articles and web sites, FactSet identifies supply chains of listed and non-listed firms with their major partners. Table 1 shows the number of all firms (for reference) and listed firms (our sample) in the FactSet data and their supply-chain links by year. The number of firms covered in the FactSet data increased rapidly over time, because the dataset has been constructed more recently and thus ignores firms in the distant past.
| 2003 | 2004 | 2005 | 2006 | 2007 | 2008 | 2009 | |
| Firms across the world (total) | |||||||
| Number of firms | 9,283 | 9,292 | 9,804 | 10,058 | 10,342 | 10,197 | 10,802 |
| Number of links | 38,594 | 40,749 | 45,172 | 38,866 | 34,894 | 33,136 | 28,640 |
| firms across the world (listed) | |||||||
| Number of firms | 633 | 726 | 783 | 797 | 745 | 760 | 1,047 |
| Number of links | 1,413 | 1,685 | 1,914 | 1,723 | 1,386 | 1,364 | 1,687 |
| 2010 | 2011 | 2012 | 2013 | 2014 | 2015 | 2016 | |
| Firms across the world (total) | |||||||
| Number of firms | 17,729 | 24,671 | 28,244 | 31,090 | 40,146 | 53,114 | 64,968 |
| Number of links | 42,971 | 61,675 | 72,671 | 82,673 | 101,748 | 133,134 | 167,772 |
| firms across the world (listed) | |||||||
| Number of firms | 2,153 | 3,842 | 4,798 | 5,779 | 8,418 | 10,729 | 12,728 |
| Number of links | 4,179 | 8,782 | 12,073 | 15,485 | 22,473 | 31,537 | 42,201 |
Finally, to identify domestic supply chains within Japan, we employ data collected by Tokyo Shoko Research (TSR) for the period 2008-2016, particularly its Company Linkage Database that covers supply chains of more than one million firms in Japan. Supply chains are captured by annual firm-level surveys by TSR that request each firm to report up to 23 suppliers and client firms of each firm. Apparently, many firms are linked with more than 23 suppliers and clients, and thus their links cannot be fully identified by their own responses. However, these links can be mostly identified by responses by their suppliers and clients. Accordingly, the maximum number of suppliers and clients is *** and ***, respectively, in 2016, implying that our data cover supply chains of Japanese firms to a great extent. The number of firms and their supply-chain links and the corresponding number for listed firms from 2008 to 2016 are presented in Table 2.
| 2008 | 2009 | 2010 | 2011 | 2012 | 2013 | 2014 | 2015 | 2016 | |
| Japan (total) | |||||||||
| Number of firms | 1,018,570 | 1,061,051 | 1,107,760 | 1,135,200 | 1,156,205 | 1,182,729 | 1,194,034 | 1,200,162 | 1,247,939 |
| Number of links | 4,443,677 | 4,644,977 | 4,810,967 | 4,925,307 | 4,995,612 | 5,090,119 | 5,157,181 | 5,209,604 | 5,491,417 |
| Japan (listed firms) | |||||||||
| Number of firms | 3,554 | 3,456 | 3,403 | 3,308 | 3,245 | 3,215 | 3,204 | 3,215 | 3,205 |
| Number of links | 21,720 | 22,128 | 22,630 | 22,044 | 21,833 | 21,696 | 21,439 | 20,968 | 20,445 |
Methods
Sentiment analysis using FinBERT
Our sentiment analysis to determine the degree of positiveness and negativeness of the information about the listed firms provided in each news article utilizes a natural language processing (NLP) model called FinBERT[40]. FinBERT is an extended version of the Bidirectional Encoder Representations from Transformers (BERT) developed by Devlin et al. (2018)[39]. Specifically, FinBERT uses Reuters’ TRC2-financial that consists of 1.8 million news articles between 2008 and 2010 and fine-tunes BERT for financial sentiment classification using Financial Phrasebank that consists of 4,845 English sentences taken from financial news in the LexisNexis database. From each news article, FinBERT generates softmax outputs for three labels, i.e., positive, neutral, and negative, which indicate the probabilities of the three about the firms mentioned in the article. Before conducting the sentiment analysis using FinBERT, we clean news articles by deleting URLs, line breaks, ISINs, and fixed phrases at the end.
The distribution of the probabilities of positive, neutral, and negative sentiments generated by FinBERT from news articles about firms across the world and Japanese firms are provided in Panels (A) and (B) of Figure 3, respectively. Both panels indicate that the probability of neutral sentiment is high for many news articles followed by that of negative sentiment while news articles with a large probability of positive sentiment are fewer.
From the probabilities of the three types of sentiment, we construct an index of positive sentiment by the ratio of the probability of positive sentiment to the sum of the probabilities of positive and negative sentiment, ignoring the probability of neutral sentiment. For example, when the probabilities of positive, neutral, and negative sentiment are 0.3, 0.5, and 0.2, respectively, our positive sentiment index is . We ignore the probability of neutral sentiment to avoid the ambiguity of sentiment in news articles. Figure 4 demonstrates the distribution of this positive sentiment index for firms across the world and Japanese firms. Because we avoid the ambiguity of the sentiment, the distribution of the index is skewed to both lower (0) and upper (1) limits although we observe a hike around the value of 0.25. In other words, our positive sentiment index distinguishes between positive and negative sentiment more clearly than the probabilities of the three directly generated by FinBERT and highlights positive sentiment in news articles.
Regression analysis
Direct effect
To estimate the effect of positive sentiment in news articles about particular firms on their own stock prices, we apply the positive sentiment index constructed above to regression analysis. Specifically, we focus on firms mentioned by any news article and examine whether the change rate of their stock prices in a time window are affected by the level of positiveness of the article after the article is disclosed, compared with its change rate before the disclosure. Suppose that firm listed in stock market is mentioned by news article on day . Setting the time window at , such as 2, 5, and 30 days, we define the average daily percentage change of the stock price of firm before and after news article disclosed on day in time window , or and , respectively, by
where is the average stock price between day and . For example, is the average stock price of firm for days between day and . Therefore, the left-hand side of equations (1) and (2) indicates the daily percentage change of the average stock price in time window in the pre- and post-news period, respectively. We take the average of stock prices over the time window to reduce their fluctuations. We exclude Saturdays and Sundays from our daily observations because stock prices are not available on weekends. Accordingly, for example, when news article about firm is disclosed on Thursday, the 10th of June and the time window is set at 3 days, the pre-news change rate is defined as the change rate from the average stock price on June 3rd (Thursday), 4th (Friday), and 7th (Monday), skipping 5th (Saturday) and 6th (Sunday), to the average on 8th (Tuesday)-10th (Thursday), as shown in Figure 5. Similarly, the post-news change is defined as the change from the average on June 7th-9th to that on June 10th, 11th (Friday), and 14th (Monday).
Thus, our sample is restricted to firm-day observations before or after news articles about the firms are disclosed. In other words, the number of observations in our sample is the cumulative number of firms mentioned in news articles during the period examined multiplied by 2. Using the pre- and post-news observations, we estimate the following estimation equation:
where and are indicator variables that take a value of one if is the percentage change of the stock price of firm before and after news article is disclosed on day , respectively, or . represents the level of positive sentiment in news article generated by FinBERT (Data section). is the percentage change of the Refinitiv index for each stock market that controls for the overall time trend of stock prices in the market, and is sector dummies. Sectors are defined by the North American Industry Classification System at the sub-sector level for the sample of firms across the world and by the Japan Standard Industry Classification at the two-digit level for the sample of Japanese firms.
Accordingly, indicates the average effect of the positiveness of a news article on the average change rate of the stock price of the firm mentioned in the article after it is disclosed in the time window of days, controlling for the overall trend in the market and sector-specific unobservable factors. By contrast, indicates any prior “effect” of the news article on the change rate of the stock price, including the possible effect of the information in the news article that diffuses in the market not through the news but through other informal channels before the disclosure of the news. In addition, may include the endogeneity bias reflecting the possibility that the growth rate of the stock price of a firm that is positively mentioned by any news article is intrinsically higher than that of a firm negatively mentioned. Therefore, the net effect of the disclosure of the information by the news article on the change rate of the stock price in time window should be given by [42].
We particularly estimate equation (3) for to examine both the short- and long-run effect of positive news articles on stock prices. We hypothesize that a positive news article would accelerate the change rate of the stock price of the firm mentioned in the article in the short-run, while the effect would become smaller and converge to 0 in the longer run. Note that equation (3) for cannot be estimated because the change rates of the stock price in the pre- and post-news periods are the same, according to equations (1) and (2).
It should be emphasized that this analysis is comparing firms positively mentioned in news articles with those negatively mentioned, rather than comparing firms positively mentioned and those not mentioned by news articles. We focus on the sub-sample of firms mentioned by news articles and ignore those not mentioned, because the two types of firms are most likely to be intrinsically different from each other to a great extent so that comparing the two may be associated with large endogeneity biases.
Indirect effect through supply chains
We further examine the effect of positive sentiment about firms in news articles on their suppliers and customers, using similar frameworks. Specifically, when we focus on the upstream effect of news articles about firms on their suppliers, in equation (3) is replaced with the probability of positive sentiment in the news article that mentions not firm but any of its customer firms. Alternatively, when we examine the downstream effect of customers’ news on their customers, is replace with the positive index of any of suppliers of firm . When a firm mentioned by a news article has more than one supplier or customer, as shown in Tables 1 and 2, we include all the suppliers and customers in our sample. In this analysis, we compare stock prices of suppliers or customers of firms mentioned positively in news articles with those of supplier or customers of firms mentioned negatively, as in our analysis on the direct effect on own stock prices.
We hypothesize that the effect of news articles on stock prices diffuses through supply chains. Precisely, a positive news article about a customer firm would accelerate the growth in the stock price of its suppliers, in addition to its own stock price, because the market predicts an increase in demand for products of the suppliers associated with higher performance in the customer firm. Also, positive news about a supplier would promote growth of its customers because of possible productivity growth of the supplier. As in the case of the direct effect, the indirect effect on suppliers and customers would be smaller in the long run than in the short run.
Results and Discussion
Direct effect of news articles on listed firms across the world
We start with estimating how the positiveness of news articles about firms affect percentage changes in their stock prices before and after the disclosure of news, applying equation (3) to the sample of publicly listed firms across the world mentioned by any news article. Results are presented in Table 3 and Figure 6.
| Dependent Variable: | Daily percentage change of the stock price of the firm mentioned in each news article in the pre- or post-news period | ||||||
| Model: | (1) | (2) | (3) | (4) | (5) | (6) | (7) |
| Time window (days) | 2 | 3 | 4 | 5 | 30 | 180 | 365 |
| Market-specific trend | Yes | Yes | Yes | Yes | Yes | Yes | Yes |
| Sector fixed effects | Yes | Yes | Yes | Yes | Yes | Yes | Yes |
| value | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 |
| Number of observations | 9,286,677 | 9,334,585 | 9,354,107 | 9,369,523 | 9,402,470 | 9,381,512 | 9,314,442 |
In column (1) of Table 3, the estimated is 0.419 and highly significant, indicating that the level of the positiveness in a news article about a firm is positively correlated with the percentage change in the firm’s stock price for 2 days before the disclosure of the news. Specifically, because the positiveness index, in equation (3), ranges from 0 to 1 (Figure 4), the value of the coefficient means that the change rate of the stock price of a firm mentioned in a news article most positively () is 0.419% points higher than that of a firm mentioned most negatively () before the disclosure of the news. There are two possibilities for this positive coefficient. First, positive and negative information included in news articles diffuses to stock markets partially through informal channels before the formal disclosure of news, generating the pre-news effect. Second, stock prices of firms positively mentioned by news articles grow faster intrinsically than those of firms negatively mentioned, leading to endogeneity biases.
Column (1) of Table 3 further finds that the estimated is 0.505, indicating that the stock price of a firm mentioned in a news article most positively grows faster by 0.505% points than that of a firm mentioned most negatively after the news disclosure. The second and third rows from the bottom show and the value from a test for the null hypothesis that the difference is 0. In column (1), the value is close to 0, and thus we reject the null hypothesis. This result suggests that although stock prices of firms mentioned by news positively grow faster than those of firms mentioned negatively even before the disclosure of news, positive news raises stock prices of former firms more after its disclosure.
These results in column (1), or for the time window of 2 days, generally hold in columns (2)-(5) for time windows of 3-5 and 30 days, while the post- and pre-news effects and the difference between the two are smaller as the time window becomes wider. By contrast, the difference between the post- and pre-news effects is negative when the time window is 180 and 365 days, or in the long run, although the difference is quite small in size. This result implies that the intrinsic difference between firms mentioned positively and negatively by news articles is negligible. Therefore, the positive and significant difference between the post- and pre-news effects in the short run (time windoes from 2 to 30 days) is mostly due to information diffusion through informal channels before the disclosure of news.
Figure 6 graphically demonstrates the argument above. The left half of the figure where the value on the horizontal axis is negative shows the point estimate of the pre-news effect for different time windows, whereas the right half shows the post-news effect. Although the confidence intervals are added in the figure, they are too small to be recognized. The estimated coefficients show an inverted-U shape where the coefficients are close to 0 on the left and right edges. This shape suggests the followings. The change rate of the stock price of firms positively mentioned by news articles is not intrinsically different from that of firms negatively mentioned long before the disclosure of the news. However, the change rate of the stock price of positively-mentioned firms started to increase compared with negatively-mentioned firms 30 days before the disclosure and reaches the peak 2 days after the disclosure. Then, the difference between the change rates of the stock prices of positively- and negatively-mentioned firms becomes smaller over time and converges to almost 0 180 days after the disclosure.
Indirect effect on suppliers through global supply chains
Next, we examine whether positive news about a firm affects stock prices of the firm’s suppliers and customers through global supply chains by estimating equation 3 where the positiveness index of news article about firm on day , , is replaced with the index for any supplier or customer of firm .
Figure 7 is the graphical presentation of the results for the effect of news about firms on their suppliers, whereas Table 4 particularly shows the difference between the post- and pre-news effects and its values. The results shown in the figure and table are similar to the results for the direct effect on own firms’ stock prices (Figure 6 and Table 3). The post- and pre-news effects are positive and significant, and their difference is positive and significant when the time window is narrow. These results suggest that positive news about firms raises stock prices of the firms’ suppliers before the disclosure of the news through informal information diffusion and raises them more after the disclosure.
| Time window (days) | 2 | 3 | 4 | 5 | 30 | 180 | 365 |
| value | 0.00 | 0.00 | 0.23 | 0.00 | 0.00 | 0.03 | 0.00 |
However, there are two notable differences between the direct effect on own firms and the indirect effect on suppliers. First, the post- and pre-news effect on suppliers and their difference is substantially smaller than the direct effects and their difference. For example, the pre-news effect on suppliers for the time window of 2 days is 0.0238 (Figure 7), while the corresponding direct effect is 0.419 (Figure 6), more than 10 times larger. Similarly, the net effect of the disclosure of news, represented by and shown in Table 4, is 0.0033 when the time window is 2 days, while the corresponding value for the direct effect is 0.086 (Table 3). Therefore, we conclude that although positive news articles about firms affect stock prices of their suppliers, the effect on suppliers is less than 10% of the direct effect on own firms. Second, the difference between the post- and pre-news effects on suppliers is positive and significant only when the time window is 2 or 3 days, although the significantly positive difference between the post- and pre-news direct effects is observed in a longer run, up to the time window of 30 days. This implies that the net effect of the disclosure of news on suppliers lasts for only several days, a very short run, after the disclosure.
Indirect effect on customers through global supply chains
Similarly, we examine the effect on customer firms and show the results in Figure 8 and Table 5. In this analysis, although the pre- and post-news effects on customers show an inverted-U shape (panel [A] of Figure 8), as in the previous results, we do not find any positive and significant difference between the post- and pre-news effects (the upper part of Table 5). Therefore, in addition to the benchmark index of positiveness of each news article using the probabilities of positive and negative sentiment from the FinBERT (Figure 4), we employ an alternative positiveness index based on only the probability of positive sentiment shown by pink bars in Figure 3. As is clear from the comparison between Figures 3 and 4, the alternative index tends to evaluate the positiveness of news articles less than the benchmark index. Thus, using the alternative index, we can examine the effect of highly positive news. Then, we find a positive and significant difference between the post- and pre-news effects, or the net effect of the disclosure of highly positive news articles, on customers for the time windows of 2 and 3 days, as shown in the lower rows of Table 5. As in the case of the effect on suppliers, the net effect on customers is small and lasts only for several days.
| 2 | 3 | 4 | 5 | 30 | 180 | 365 | |
| Benchmark positiveness index using the positive and negative probabilities of news articles | |||||||
| 0.0008 | -0.0007 | ||||||
| value | 0.21 | 0.14 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 |
| Alternative positiveness index using only the positive probability of news articles | |||||||
| value | 0.00 | 0.00 | 0.44 | 0.00 | 0.00 | 0.00 | 0.00 |
Effect on Japanese firms
We further use another sample that focuses on Japanese listed firms. The results for the direct effect of positive news articles on Japanese firms’ own stock prices shown in Figure 9 are qualitatively the same as and quantitatively similar to those using the sample of firms across the world shown in Figure 6.
In addition, we examine the indirect effect of news articles about firms on stock prices of their suppliers and customers. A notable difference between this analysis on Japanese firms and the previous one on firms across the world is that more detailed domestic supply chains are identified based on firm-level surveys in the former than in the latter that mostly relies on financial statements (Data Section). Therefore, the average number of supply chain links for each Japanese firm in 2016 is 6.4, while that for each firm in the world is 3.3 (Tables 1 and 2).
The results are presented in Figure 10 and Table 6. Panels (A) and (B) of Figure 10 for the effect on suppliers and customers, respectively, show several characteristics similar to each other but different from the previous results. First, although the pre- and post-news effect on stock prices of both suppliers and customers are mostly positive and significantly different from 0, as in the previous results, the pre-news effect is negative when the time window is 180 and 365 days, unlike the previous results. The negative pre-news effect in the long run implies that stock prices of firms positively mentioned in news do not necessarily grow faster long time before the news. In other words, the effect of news is less likely to be overestimated due to selection of potentially fast-growing firms.
| Time window (days) | 2 | 3 | 4 | 5 | 30 | 180 | 365 |
| Effect on suppliers | |||||||
| value | 0.21 | 0.14 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 |
| Effect on customers | |||||||
| value | 0.00 | 0.00 | 0.44 | 0.00 | 0.00 | 0.00 | 0.00 |
Second, the pre-news effect is positive and significant, implying diffusion of information before the disclosure of news that is consistent with the previous results. However, the difference between the post- and pre-news effects that can be interpreted as the net effect of the disclosure of news is negative in the short run, or in the time window of 2-5 days and turns to be positive in the long run (Table 6). This result is completely opposite to the previous ones using data for firms across countries (Tables 4 and 5). In other words, positive information about a firm diffuses through informal channels and raises the change rate of stock prices of the firm’s suppliers and customers before the information is disclosed in a news article. However, the disclosure of the information by a news article lowers the change rate of stock prices compared with that in the pre-news period, although the post-news effect is still positive. One possible reason for this is that investors who can receive information about firms through informal channels are more knowledgeable and recognize diffusion of performance through supply chains more accurately than other investors who rely on formal news. Therefore, once the knowledgeable investors receive positive information about a firm before the disclosure by news, they invest in the firm’s suppliers and customers. However, other investors react less to positive news about firms’ suppliers and customers.
Finally, the pre- and post-news effects on suppliers and clients found in the sample of Japanese firms are larger than those in the sample of firms across the world. This finding, combined with the previous interpretation about knowledgeable investors, suggests that supply chain ties between suppliers and customers in Japan, known as keiretsu, are often stronger than in other countries, because they are often associated with shareholding relationships, information exchanges, technical assistance and research collaboration.[43, 44] Therefore, diffusion of performance through supply chains may be more prominent in Japan than in others, and this possibility is recognized by at least knowledgeable investors in Japan.
Conclusion
This study examines whether positive news about firms increases their stock prices and, moreover, whether it increases stock prices of the firms’ suppliers and customers, using a large sample of publicly listed firms across the world and another of Japanese listed firms. The level of positiveness of each news article is determined by FinBERT, a natural language processing model fine-tuned specifically for financial information. Supply chains of firms across the world are identified mostly by financial statements, while those of Japanese firms are taken from large-scale firm-level surveys.
We find that positive news increases the change rate of stock prices of firms mentioned in the news before its disclosure, most likely because of diffusion of information through informal channels. Positive news also raises stock prices of the firms’ suppliers and customers before its disclosure, confirming propagation of market values through supply chains. In addition, we generally find a larger post-news effect on stock prices of the mentioned firms and their suppliers and customers than the pre-news effect. The positive difference between the post- and pre-news effects can be considered as the net effect of the disclosure of positive news, controlling for informal information diffusion. However, the post-news effect on suppliers and customers in Japan is smaller than the pre-news effect, a result opposite to those from firms across the world. This notable result is possibly because supply chain links of Japanese firms are stronger than global supply chains while such knowledge is restricted to selected investors.
Acknowledgements
The authors thank Yuito Nishi for his technical contribution.
Author contributions statement
H.I conceived the study. H.I and Y.T conducted the analyses and wrote and reviewed the manuscript.